### v2

# Collapsing in polygonal dynamics

Abstract
Polygonal dynamics is a family of dynamical systems containing many studied systems, like the famous pentagram map. Similar collapsing phenomena seem to occur in most of these systems. We give a unifying, general definition of polygonal dynamics, and conjecture that a generic orbit collapses towards a predictable point. We manage to prove it in some setting. For the special case of “closed polygons”, we show as a corollary that the collapse point depends algebraicly on the vertices of the starting polygon, using tools called scaling symmetry and infinitesimal monodromy. This generalizes previous results about the pentagram map. Then, we investigate the case of polygonal dynamics in $\mathbb{P}^{1}$ for which we give an explicit polynomial equation satisfied by the collapse point. Based on previous works, we define a new dynamical system, the “staircase” cross-ratio dynamics, for which we study particular configurations.
1 Introduction
The point of this paper is to study a family of dynamical systems, called polygonal dynamics. It stems from observing similar behaviours when numerically simulating them. Let’s start with a definition.
Definition 1.1.
Let $n,\penalty 10000\ d$ be positive integers. A twisted $\bm{n}$-gon $P$ is the data of $(p_{1},\dots,p_{n})$ and $M$, where the $p_{i}\in\mathbb{P}^{d}$ are the vertices and $M\in\mathbb{P}\mathrm{GL}_{d+1}$ is the monodromy. The vertices are to be thought as an infinite sequence satisfying the condition:
$p_{i+n}=Mp_{i}\quad\forall i\in\mathbb{Z}.$
$P$ is said to be closed if $M=\mathrm{Id}$. The set of twisted and closed $n$-gons are respectively denoted by $\tilde{\mathcal{T}}_{n}$ and $\tilde{\mathcal{P}}_{n}$.
A polygonal transformation is a birational map $T:\tilde{\mathcal{T}}_{n}\dashrightarrow\tilde{\mathcal{T}}_{n}$ which preserves the monodromy and commutes with the following action of $\mathbb{P}\mathrm{GL}_{d+1}$:
$A\cdot P=((Ap_{1},\dots,Ap_{n}),AMA^{-1}),\quad A\in\mathbb{P}\mathrm{GL}_{d+1}.$
The class of $P$ is denoted by $[P]$. The action of $\mathbb{P}\mathrm{GL}_{d+1}$ allows to define the quotient spaces [^2] Sometimes taking a quotient is more tedious, and we need to use geometric invariant theory (GIT). See [1, §3.1] for details. $\mathcal{T}_{n}$ and $\mathcal{P}_{n}$, also called moduli spaces.
Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$, it also yields a dynamic over the quotient spaces. As we will see later, it is often very different from the one on the initial space.
Remark 1.
We denote the projective spaces by $\mathbb{P}^{d}$ and their automorphism groups by $\mathbb{P}\mathrm{GL}_{d+1}$ without specifying over which field $k$ they are defined. This is because we see them from the algebraic geometry standpoint, as schemes (like it’s done in [1]). Similarly, the group scheme $\mathbb{G}_{m}$ refers to the group of invertible elements $k^{*}$. However, no algebraic geometry background will be needed here, and the unfamiliar reader may consider the statements over a fixed field, like $\mathbb{C}$.
Remark 2.
This definition can be generalized by replacing $\mathbb{P}^{d}$ by a topological space $X$ and $\mathbb{P}\mathrm{GL}_{d+1}$ by a group $G$ acting by homeomorphism (see [2, §2]). For instance, the polygon recutting defined by Adler [3] in 1993 takes place in $X=\mathbb{R}^{2}$ with $G=\mathrm{Isom}^{+}(\mathbb{R}^{2})$. See table LABEL:tab:examples_polygonal_dynamics for a review of existing polygonal dynamics.
Since our results only concerns polygonal dynamics in $\mathbb{P}^{d}$, we will set the rest aside for now.
Figure 1: The convex pentagon $P$, lying on the real plane, is sent to $P^{\prime}$ by the pentagram map.
Let’s start with the most famous polygonal dynamic: the pentagram map, defined by Schwartz [4] in 1992. In his original paper, he defines the map $T$ acting on the set of strictly convex closed $n$-gons (with $n\geq 5$, hence the name) in $\mathbb{P}^{2}(\mathbb{R})$. [^3] He generalised it to twisted polygons [5] in 2007. A polygon $P=(p_{1},\dots,p_{n})$ is sent to another one by taking the intersections of the small diagonals $\overline{p_{i-1}p_{i+1}}$ and $\overline{p_{i}p_{i+2}}$, where the indices are understood modulo $n$ (see figure 1 for an example). Schwartz proved that for these real and strictly convex polygons, the sequence of iterates collapses exponentially fast towards a point. However, the meaning of this point, or even the existence of a formula to express it from the original vertices of $P$, remained unknown for 26 years. This collapse point, which can be thought as a kind of “projective center of mass” of the polygon $P$, is quite intriguing.
In 2018, Glick [6] gave a formula for the collapse point. To do this, he defines a linear operator associated to a convex polygon $P$, now called Glick’s operator, which is invariant by the dynamic and sends $P$ to its interior. He concludes that the collapse point must be a fixed point of the operator. The formula is quite simple, involving only a cubic extension of the field generated over $\mathbb{Q}$ by the coordinates of the vertices of $P$. However, the proof is quite computational and doesn’t shed much light on the meaning of the point.
In 2020, Aboud and Izosimov [7] reinterpret this point as being an eigenline of a certain $3\times 3$ matrix $M^{\prime}$ associated to the polygon $P$, called its infinitesimal monodromy. This matrix can be in some way identified with Glick’s operator. To do this, they “open” the closed polygon into a family of twisted ones through a tool called the scaling symmetry.
Definition 1.2.
A scaling symmetry is a rational group action $\rho:\mathbb{G}_{m}\times\mathcal{T}_{n}\to\mathcal{T}_{n}$ which commutes with the dynamic of $T$.
But some questions remain unanswered:
  1. 1.
    Which of the three eigenlines gives the collapse point, for what reason, and what is the meaning of the other ones ?
  2. 2.
    In [8, §9.2], Schwartz notices that for a family of transformations $\Delta_{k}$ (for which $\Delta_{1}$ is the pentagram map) and for certain polygons (called $k$-birds), there is a collapse behaviour like for the pentagram map with convex polygons. Moreover, the coordinates of collapsing are again given by a cousin of Glick’s formula. To quote him: “I wonder if this means that the collapse point exists for all starting points of the pentagram map. Even if the iterations go completely crazy under the map, perhaps they still collapse to the point predicted by Glick’s operator. The idea of a completely general collapse point has always seemed absurd to me, but maybe it is not. Nobody knows.
  3. 3.
    In [9, §7.2], Arnold and Arreche notice a similarity between the infinitesimal monodromy of the pentagram map and a so-called “hyperbolic barycenter” defined for the flat cross-ratio dynamics (a dynamical system for polygons on $\mathbb{P}^{1}(\mathbb{R})$, see [10, §6.2.1]). But the link is unclear.
In this paper, we answer these questions by considering the general setting of any polygonal dynamic with some additional hypotheses. From numerical experimentation, we conjecture that the collapsing generically [^4] By generic, we mean for a dense open subset of polygons, where the topology is given by the metric. occurs. This is reminiscent of an idea of Schwartz in his seminal paper about the pentagram map.
Conjecture 1.3 – 3.2.
Let $k$ be an algebraicly closed, complete field for the metric coming from a non-trivial valuation. Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$, such that the dynamic on the moduli space is discretely integrable [^5] See definition 2.4..
Then for a generic choice of $n$-gon $P=((p_{1},\dots,p_{n}),M)$, there exists a matrix $\tilde{M}$ commuting with $M$, such that the vertices of $P$ collapse under the iteration of $T$ in the past/future towards $q_{-}$ and $q_{+}$, which are the repelling/attracting fixed points of $\tilde{M}$. Moreover, $q_{-}$ and $q_{+}$ depend continuously on $P$.
This brings an answer to question 1. We prove it in some specific configurations, for which we can exhibit some examples. It is important to remind it hold generically, and not for all polygons, since we find some non-generic counter-examples in subsection 6.2.
Theorem 1.4 – $3.$
Conjecture 3.2 holds for polygons whose dynamic is periodic on the moduli space.
Now, we want a formula for the collapse points. From lemma 3.1, we obtain that they have to be fixed points of the monodromy $M$ of the polygon $P$. However, we were originally interested in closed polygons only, and this is useless when $M=\mathrm{Id}$.
To resolve this issue, we use the scaling symmetry to define an infinitesimal monodromy $M^{\prime}$ (see lemma 2.1 which generalizes [7, Prop. 3.2]). This allows to retrieve the “true fixed points of the identity”, from which we obtain corollary 3.4.
Corollary 1.5 – 3.4.
Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$ where $k$ is a complete non-trivially valued field, such that the dynamic is integrable on the moduli space and admits a scaling symmetry. Let $P$ be a generic closed polygon and $M^{\prime}$ be its infinitesimal monodromy.
Then the vertices of $P$ collapse under the iteration of $T$ in the past/future toward fixed point of $M^{\prime}$.
An other corollary is the generalization of [6, Thm 1.1]. Since the infinitesimal monodromy is a $(d+1)\times(d+1)$ matrix, finding its eigenlines only requires solving a linear system.
Corollary 1.6 – 3.5.
Let $P$ be a closed $n$-gon in $\mathbb{P}^{d}$ and $(p_{i}=[p_{i}^{(0)}:\dots:p_{i}^{(d)}])_{i=1,\dots,n}$ be its vertices. Let $q=[q^{(0)}:\dots:q^{(d)}]$ be a collapse point predicted by corollary 3.4. Then $q^{(0)},\dots,q^{(d)}$ are contained in a field extension of $\mathbb{Q}\big((p_{i}^{(j)})_{i=1,\dots,n}^{j=0,\dots,d}\big)$ of degree at most $d+1$.
This provides answers to questions 2 and 3 raised earlier.
Indeed, the pentagram map and its generalizations admit scaling symmetries and thus infinitesimal monodromies, so this answers Schwartz’s question about the reappearance of Glick’s formula.
This also explains the similarities observed by Arnold and Arreche between Glick’s formula for the pentagram map and the “hyperbolic barycenter” for cross-ratio dynamics. In fact, for any polygonal dynamic in $\mathbb{P}^{1}$ admitting a scaling symmetry, we give an explicit formula for the infinitesimal monodromy. Moreover, we give a degree $2$ polynomial equation that the collapse points have to verify, thus making an effective version of corollary 3.5.
Theorem 1.7 – 4.2.
Suppose to have a polygonal dynamic on $\mathbb{P}^{1}$ admitting a scaling symmetry. Let $P$ be a closed $n$-gon, and $c_{i}(z)$ be the cross-ratio coordinates of $P_{z}$. Then the infinitesimal monodromy of $P$ is
$M^{\prime}=\begin{pmatrix}n/2+J&-K\\ I&n/2-J\end{pmatrix}$,
where
$I=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}},\quad J=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}+p_{i+1}}{p_{i}-p_{i+1}},\quad K=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}p_{i+1}}{p_{i}-p_{i+1}}.$
Furthermore, its eigenvectors $(X,Y)^{T}$ satisfy the polynomial equation
$\chi_{P}(X,Y):=IX^{2}-2JXY+KY^{2}=0.$
This theorem also retrieves the invariant quantities $I,J,K,$ that appeared as particular cases in [10, §6.2] and [9]. Ideally, we could even retrieve an invariant pre-symplectic two form which keeps reappearing in each polygonal dynamics in $\mathbb{P}^{1}$, but we didn’t manage. The answer may come from a deeper interpretation in term of $\lambda$-lengths (see [10, §5.3]).
Finding the scaling symmetry of a system can be quite complicated. Once it’s done, we can apply theorem 4.2 to predict the collapse points. It is also the main tool to prove algebraic integrability (see [1] for an example on any algebraicly closed field). In this paper, we compute the scaling symmetry of the leapfrog map (defined in [11, §5.2]), the “flat” cross-ratio dynamic (studied in [10]), and the “staircase” cross-ratio dynamic. The last system originates from the “staircase” geometry in the theory of continuation of discrete holomorphic functions (see [12, §6] and references therein). However, it is more general than the initial problem since we allow to do the “flip” continuation on any vertices, regardless of any geometry. It makes up for a whole group of transformations, which we prove to be the affine symmetric group (see [13]). This group appeared in another polygonal dynamical system, the polygon recutting [3], which takes place in $\mathbb{R}^{2}$.
Outline of the paper: In section 2, we give some basic definitions and properties, and a review of known polygonal dynamics. In section 3 we formulate the conjecture 3.2 about generic collapsing. Then we focus on the closed polygons, for which corollary 3.4 predicts the collapse points. Section 4 is focused on polygonal dynamics in $\mathbb{P}^{1}$, for which theorem 4.2 gives an effective version of the corollary 3.4. We apply it in section 5 to some previously defined dynamics, and in section 6 to a new one, the staircase cross-ratio dynamics. We conclude by studying particular orbits of this system, which provides hints in favor of our conjecture.
2 Basic properties and examples
We are interested in polygonal dynamic $T$ taking place in $\mathbb{P}^{d}$ and admitting a scaling symmetry (see definitions 1.1 and 1.2). However, many reasonings would apply to more general cases (see [2, §2]). At the end of this section, table LABEL:tab:examples_polygonal_dynamics presents some known polygonal dynamics and their properties.
Let’s start by defining the main tool we’ll use for studying the dynamic on closed polygons.
Definition/Proposition 2.1.
Let $\rho$ be a scaling symmetry for a certain polygonal dynamic over $\mathbb{P}^{d}(k)$, where $k$ is a complete field.
Let $P]\in\mathcal{P}_{n}$ and $P]_{z}:=\rho_{z}\cdot[P]$. Up to a choice of a family of projective transformations, we can choose a smooth lift of the deformation
$P_{z}=((p_{1}(z),\dots,p_{n}(z)),M_{z})\in\tilde{\mathcal{T}}_{n}.$
Define the infinitesimal monodromy to be:
$M^{\prime}:=\left.\frac{dM_{z}}{dz}\right|_{z=1}.$
It is independent of the choice of lift and it is invariant under $T$.
Proof.
It is exactly same proof as in [7, Prop 3.2], which treated the case of $\mathbb{P}^{2}(\mathbb{C)}$. The deformation by the scaling symmetry takes place on $\mathcal{T}_{n}$, but we can still take a smooth lift on $\tilde{\mathcal{T}}_{n}$ which will commute with the dynamic. One example [^6] Which is the one we use in section 4. of such a lift is the family of representants $(P_{z})$ for which the first $d+1$ vertices $p_{1}(z),p_{2}(z),\dots,p_{d+1}(z)$ are fixed on
$1:\dots:1],[1:0:\dots:0],\dots,[0:\dots:0:1].$
Because of [14, Prop. 4.5.10], this is the unique family of representant of $([P]_{z})$ satisfying this property.
Now, let $P_{z},\penalty 10000\ \tilde{P}_{z}$ be two lifts such that $P_{1}=\tilde{P}_{1}$, and $\phi_{z}\in\mathbb{P}\mathrm{GL}_{d+1}$ be a smooth family such that $\tilde{P}_{z}=\phi_{z}\cdot P_{z}$. This forces $\phi_{1}=\mathrm{Id}$ and $\tilde{M}_{z}=\phi_{z}M_{z}\phi_{z}^{-1}$. By differentiating, which we can do since $k$ is a complete metric space, we get:
$\left.\frac{d\tilde{M}_{z}}{dz}\right|_{z=1}=\left.\frac{dM_{z}}{dz}\right|_{z=1}+\left[\left.\frac{d\phi_{z}}{dz}\right|_{z=1},M_{1}\right].$
Since $P_{1}$ is a closed polygon, we have $M_{1}=\mathrm{Id}$ and so the commutator vanishes. Hence infinitesimal monodromy is well defined. [^7] As it was remarked in [7], the infinitesimal monodromy is not well defined for twisted polygons, since the only central element of $\mathbb{P}\mathrm{GL}_{d+1}$ is $\mathrm{Id}$.
The infinitesimal monodromy is defined from the family of monodromies $(M_{z})_{z\in Z}$ coming from the scaling symmetry. But $T$ preserves the monodromy and commutes with the scaling symmetry, so it also preserves the infinitesimal monodromy. ∎
The rest of this section are tools for the collapse conjecture 3.2. Next lemma says that if $k$ is an algebraicly closed complete valued field, for a non-trivial valuation, then iterating a generic element of $\mathbb{P}\mathrm{GL}_{d+1}(k)$ yields a North-South dynamic.
Lemma 2.2.
Let $k$ be an algebraicly closed, complete field for the metric coming from a non-trivial valuation.
Then a generic element of $\mathbb{P}\mathrm{GL}_{d+1}(k)$ has one attractive and one repelling fixed point.
Proof.
Let $M\in\mathrm{GL}_{d+1}(k)$. The set of diagonalisable matrices of $\mathrm{GL}_{d+1}(k)$ is Zariski dense; indeed, $M$ is diagonalisable if its characteristic polynomial $\chi_{M}(X)$ have distinct roots. This happens if its discriminant, which is a polynomial in the coefficients $\Delta(a_{1,1},\dots,a_{d+1,d+1})$, doesn’t vanish. Since this is not the zero polynomial, we obtain a Zariski-open set. It is not empty since it contains the identity.
So up to conjugation we can assume that:
$M=diag(\lambda_{1},\dots,\lambda_{d+1}),$
with $|\lambda_{1}|\leq\dots\leq|\lambda_{d+1}|$. Since the valuation is non-trivial, then out of genericity we can assume that the inequalities are in fact strict. In $\mathbb{P}\mathrm{GL}_{d+1}$, we can take the representant with $\lambda_{d+1}=1$. So:
$\lim_{N\to\infty}M^{N}=E_{d+1,d+1}.$
Thus $M$ has an attractive fixed point in $\mathbb{P}^{d}$, being $0:\dots:0:1]$ (the origin).
We can do similarly for $M^{-1}$ and take a representant with $\lambda_{1}=1$. This yields:
$\lim_{N\to\infty}M^{-N}=E_{1,1},$
and the repelling fixed point of $M$ is $1:0:\dots:0]$ (“the” point at infinity). ∎
The following fact is classical but we’ll prove it anyway. If we apply it in $\mathbb{P}\mathrm{GL}_{d+1}$, shared eigenlines translates to shared fixed points.
Lemma 2.3.
Let $M,M^{\prime}\in\mathrm{GL}_{d+1}$ be two commuting matrices such that all the eigenvalues of $M$ are distinct. Then $M$ and $M^{\prime}$ share the same eigenlines.
Proof.
Let $x$ be such that $Mx=\lambda x$. Hence we have:
$MM^{\prime}x=M^{\prime}Mx=\lambda M^{\prime}x,$
so $M^{\prime}x$ is an eigenvector of $M$ for the eigenvalue $\lambda$. Since the eigenvalues are all distinct, $M^{\prime}x$ and $x$ lie on the same eigenline of $M$, and thus $x$ is an eigenvector of $M^{\prime}$. ∎
Let’s finish this section with a definition of discrete integrability, inspired by [1]. It aims to retrieve the essential property of the various flavours of integrability: the Liouville-Arnold one (involving symplectic forms, hamiltonians and lagrangian toris) or the algebraic one (involving Lax pairs, theta functions and abelian varieties) [^8] See [15] for a review of integrable systems.. The polygonal dynamics we consider are usually integrable on their moduli space.
Definition 2.4.
A dynamical system $(X,T)$ is said to be discretely integrable if there is a family $A$ of lagrangian tori/abelian varieties which foliate $X$ such that, on a subset of full measure $\Sigma\subset X$, there is a diffeo/birational map $\delta:\Sigma\to A$ which identifies some iterate of $T$ with a relative translation $\tau:A\to A$.
A translation can be periodic, but generically it is quasi-periodic, meaning it winds up around the torus/abelian variety in a recurrent, equidistributed way.
3 Polygonal dynamic in $\mathbb{P}^{d}$
In any of the dynamics we could simulate with SageMath over $\mathbb{C}$ and $\mathbb{Q}_{p}$ (unfortunately $\mathbb{C}_{p}$ is not implemented), we observe a collapse behaviour for the twisted and closed polygons. This is the motivation of the following section.
3.1 The collapse conjecture
A simple observation shows what are the potential collapse points. Generically, there will be only a finite number of candidates.
Lemma 3.1.
Let $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon. Suppose that under the iteration of $T$, the vertices of P all collapse towards the same value $q$. Then $q$ has to be a fixed point of $M$.
Proof.
Write $P^{(N)}:=T^{N}(P)$. Since $T$ preserves the monodromy, we have that for all $N\in\mathbb{Z}$:
$p_{i+n}^{(N)}=Mp_{i}^{(N)}.$
But since $M$ is continuous, by taking the limit we get:
$q=Mq.\qed$
Let’s formulate the conjecture. It’s inspired by a question raised by Schwartz in his original paper [4, §4]. He stated in our private communication that no progress have been achieved since 1992.
Conjecture 3.2.
Let $k$ be an algebraicly closed, complete field for the metric coming from a non-trivial valuation. Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$, such that the dynamic on the moduli space is discretely integrable.
Then for a generic choice of $n$-gon $P=((p_{1},\dots,p_{n}),M)$, there exists a matrix $\tilde{M}$ commuting with $M$, such that the vertices of $P$ collapse under the iteration of $T$ in the past/future towards $q_{-}$ and $q_{+}$, which are the repelling/attracting fixed points of $\tilde{M}$. Moreover, $q_{-}$ and $q_{+}$ depend continuously on $P$.
We only managed to prove the conjecture in the case of periodic motion on the moduli space. We know that such a periodic motion is possible for the pentagram map (see [27] and the link with Poncelet polygons) and for the staircase cross ratio dynamics (see subsection 6.2).
More generally, periodic motion over the moduli space occurs for polygons with $n$ is small enough, since their invariant tori/abelian varieties are points. As shown in [4], this happens when $n=5,6$ for the pentagram map over closed polygons. It also happens for closed $3$-gons in $\mathbb{P}^{1}$, since the moduli space is a point.
Theorem 3.3.
Conjecture 3.2 holds for polygons whose dynamic is periodic on the moduli space.
Proof.
Let $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon and $N\in\mathbb{N}^{*}$ be such that $T^{N}([P])=[P]$. In the initial space, this means that there exists $\tilde{M}\in\mathbb{P}\mathrm{GL}_{d+1}$ such that $T^{N}(P)=\tilde{M}(P)$. By lemma 2.2, we can assume that $\tilde{M}$ generically has an attractive and a repelling fixed point. Since $T$ preserves the monodromy of $P$, we obtain:
$M=\tilde{M}M\tilde{M}^{-1}.$
Since $T$ commutes with the action of $\mathbb{P}\mathrm{GL}_{d+1}$, we have:
$T^{2N}(P)=T^{N}(\tilde{M}(P))=\tilde{M}(T^{N}(P))=\tilde{M}^{2}(P).$
By recurrence (that we can also apply backwards), we get:
$k\in\mathbb{Z},\penalty 10000\ T^{kN}(P)=\tilde{M}^{k}(P).$
We also have that for all $1\leq q<N,\penalty 10000\ k\in\mathbb{{Z}}$:
$T^{kN+q}(P)=T^{q}(T^{kN}(P))=T^{q}(\tilde{M}^{k}(P))=\tilde{M}^{k}(T^{q}(P)).$
We can assume that none of the $T^{q}(P)$ contains a vertex which is fixed by $\tilde{M}$. If we denote by $q_{\pm}$ the attractive/repelling fixed points of $\tilde{M}$, then:
$\lim_{k\to\pm\infty}T^{k}(P)=((q_{\pm},\dots,q_{\pm}),M)\qed$
Now the problem is: how to prove it for quasi-periodic motion, which is the truly generic case ? Since the motion is recurrent on the moduli space, one could try to mimic the case of periodic motion to say that some subsequence of the iterate “behaves asymptotically the same” as the iteration of a projective transformation.
The meaning of “behaving asymptotically the same” is yet unclear. One strategy could be to use generalized schwarzian derivatives (see [28]), but we haven’t succeeded.
3.2 Corollaries for closed polygons
The previous conjecture combined with lemma 3.1 allows us to characterize the collapse points as being the one of the monodromy of $P$. However, the original questions were formulated for closed polygons, and $\mathrm{Id}$ fixes every point, so we cannot predict the collapse points.
If the system we’re interested in admits a scaling symmetry [^9] Which should be the case, since this is how integrability is usually obtained., we can use the infinitesimal monodromy of a polygon to retrieve the “true fixed points” of the identity. The following corollary gives us a non-trivial analogue of lemma 3.1. Because of theorem 3.3, it is true for polygons with periodic motion on the moduli space, which we already gave some examples of.
Corollary 3.4.
Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$ where $k$ is a complete non-trivially valued field, such that the dynamic is integrable on the moduli space and admits a scaling symmetry. Let $P$ be a generic closed polygon and $M^{\prime}$ be its infinitesimal monodromy.
Then the vertices of $P$ collapse under the iteration of $T$ in the past/future toward fixed point of $M^{\prime}$.
Proof.
Let $P$ be a generic closed polygon. According to conjecture 3.2, we know that the vertices of $P$ converge in the past/future towards $q_{-}$ and $q_{+}$, which are the repelling/attracting fixed points of $\tilde{M}$.
By applying the scaling symmetry, we obtain a family of polygons $P_{z}=((p_{1}(z),\dots,p_{n}(z)),M_{z})$. At least for each $z$ in a small neighborhood of $1$, we get by conjecture 3.2 that they collapse to some points $q_{-}(z)$, $q_{+}(z)$ which depend continuously on $z$. By genericity, we can assume that $M_{z}$ have distinct eigenvalues for $z\neq 1$ in this neighborhood. Since $\tilde{M}_{z}$ commutes with $M_{z}$, lemma 2.3 gives that $q_{-}(z)$, $q_{+}(z)$ are fixed points of $M_{z}$.
For $z\sim 1$ we have
$M_{z}=\mathrm{Id}+(z-1)M^{\prime}+o(z-1),$
so the fixed points of $M_{z}$ tend towards the ones of $M^{\prime}$. Since $q_{\pm}=q_{\pm}(1)$, we get the result. ∎
Remark 3.
Here are some details about its application to polygons with periodic motion on the moduli space, since the continuity part was not addressed in theorem 3.3.
The scaling symmetry commutes with the dynamic, so the dynamic of $P_{z}$ have the same periodicity on the moduli space for any $z$:
$T^{N}(P_{z})=\tilde{M}_{z}P_{z}.$
Since $\tilde{M}_{z}$ maps $(p_{1}(z),\dots,p_{d+1}(z))$ to $(p_{1}^{(N)}(z),\dots,p_{d+1}^{(N)}(z))$ and both $d+1$ tuples vary smoothly, we get by the fundamental theorem of projective geometry (see [14, Prop. 4.5.10]) that $\tilde{M}_{z}$ varies smoothly too. Hence its fixed points too. The nature of the repelling and attracting fixed points don’t change on a small enough neighborhood, so we have the result.
We also get a generalization of [6, Thm. 1.1].
Corollary 3.5.
Let $P$ be a closed $n$-gon in $\mathbb{P}^{d}$ and $(p_{i}=[p_{i}^{(0)}:\dots:p_{i}^{(d)}])_{i=1,\dots,n}$ be its vertices. Let $q=[q^{(0)}:\dots:q^{(d)}]$ be a collapse point predicted by corollary 3.4. Then $q^{(0)},\dots,q^{(d)}$ are contained in a field extension of $\mathbb{Q}\big((p_{i}^{(j)})_{i=1,\dots,n}^{j=0,\dots,d}\big)$ of degree at most $d+1$.
Proof.
Since the scaling symmetry acts by birational transformations, then the coefficients of the infinitesimal monodromy are rational functions in $p_{i}^{(j)}$. Because of corollary 3.4, a potential collapse point $q=[q^{(0)}:\dots:q^{(d)}]$ is an eigenline of this $(d+1)\times(d+1)$ matrix. To find it, one needs to solve a linear system with $(d+1)$ equations involving the coefficient of the matrix, hence the result. ∎
In the case of the pentagram map and its generalizations, we can apply the corollaries. Note that convexity is not involved. This answers the questions raised by Schwartz in [8, §9.2], at least over $\mathbb{C}$.
Corollary 3.6.
Let $k$ be an algebraicly closed complete valued field. For the pentagram map and its generalisations in $\mathbb{P}^{2}(k)$, the collapse points are fixed points of Glick’s operator.
Proof.
The generalisation of Glick’s lemma to the cousins of the pentagram maps in $\mathbb{P}^{2}$ is achieved by [8, Lemma 9.6]. Aboud and Izosimov proved in [7] that Glick’s operator coincides with the infinitesimal monodromy, and argument in fact holds on any complete field. Combined with corollary 3.4, we get the result. ∎
Remark 4.
It would be interesting to consider a similar statement in the setting of polygonal dynamics on $\mathbb{R}^{d}$ with the group $\mathrm{Isom}^{+}(\mathbb{R}^{d})$. For instance in $\mathbb{R}^{2}$, Benoist and Hulin prove in [26] that when $n=4$, the iteration of some element of the folding group is generically conjugated, on the moduli space, to a translation on an elliptic curve. This matches definition 2.4 we gave for integrability. The authors conclude that, on the initial space, the polygons drifts towards a point on the line at infinity, following some “drifting vector”. When seen on the projective plane, the fixed points of positive isometries are on the line at infinity, which is coherent with our conjecture.
4 Polygonal dynamics in $\mathbb{P}^{1}$
For any polygonal dynamic in $\mathbb{P}^{1}$, we can make a precise version of corollary 3.4 and corollary 3.5. This is presented is theorem 4.2.
We will apply it to three kinds of polygonal dynamics: the leapfrog map, the “flat” cross-ratio and the newly defined “stair-case” cross-ratio dynamics (the only ones we know). The two last ones are linked to the theory of discrete holomorphic functions (see [12, §6]) and are endowed with a discrete curvature, which is an $n$-periodic sequence $\mu=(\mu_{i})_{i\in\mathbb{Z}}$, where each $\mu_{i}\in\mathbb{G}_{m}$ should be thought to live on the edge linking $p_{i}$ and $p_{i+1}$. They are left unchanged by the action of $\mathbb{P}\mathrm{GL}_{2}$, but can be modified by the scaling symmetry.
For any $n$-gon in $\mathbb{P}^{1}$, define its cross-ratio coordinates to be:
$c_{i}:=[p_{i},p_{i+1},p_{i-1},p_{i+2}]=\frac{p_{i}-p_{i-1}}{p_{i}-p_{i+2}}\frac{p_{i+1}-p_{i+2}}{p_{i+1}-p_{i-1}},\quad i\in\mathbb{Z}.$
They are $n$-periodic and invariant under projective transformations.
Remark 5.
Sometimes another convention is taken for the cross-ratio, which is:
$(a,b,c,d)=\frac{a-b}{b-c}\frac{c-d}{d-a}.$
We have the relation:
[a,b,c,d]=1(a,b,c,d).[a,b,c,d]=1-(a,b,c,d).
In order to stay coherent with [10], we’ll stick to the first convention.
Here’s a lemma that retrieves the monodromy of any polygon in $\mathbb{P}^{1}$. Its proof can be found in the original paper.
Lemma 4.1 – [10, Lemma 3.2].
Let $P=((p_{1},\dots,p_{n}),M)$ be a $n$-gon in $\mathbb{P}^{1}$ and $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$ be its cross-ratio coordinates. Up to conjugation, suppose that $p_{0}=1,\penalty 10000\ p_{1}=\infty,\penalty 10000\ p_{2}=0$. Then in this projective chart, the monodromy matrix is:
$M=\left({\begin{array}[]{cc}0&c_{1}\\ -1&1\\ \end{array}}\right)\dots\left({\begin{array}[]{cc}0&c_{n}\\ -1&1\\ \end{array}}\right).$
When we have a scaling symmetry, this gives us a formula for the infinitesimal monodromy. Combined with corollary 3.4, this tells us what can be the collapse points for the dynamics.
Theorem 4.2.
Suppose to have a polygonal dynamic on $\mathbb{P}^{1}$ with a scaling symmetry $\rho$. Let $P$ be a closed $n$-gon, and $c_{i}(z)$ the cross-ratio coordinates of $P_{z}$. Then the infinitesimal monodromy of $P$ is
$M^{\prime}=\begin{pmatrix}n/2+J&-K\\ I&n/2-J\end{pmatrix}$,
where
$I=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}},\quad J=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}+p_{i+1}}{p_{i}-p_{i+1}},\quad K=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}p_{i+1}}{p_{i}-p_{i+1}}.$
Furthermore, its eigenvectors $(X,Y)^{T}$ satisfy the polynomial equation
$\chi_{P}(X,Y):=IX^{2}-2JXY+KY^{2}=0.$
Remark 6.
A similar matrix already appeared in [10, §6.2.1]. This answers the question asked by Arnold and Arreche at the end of [9], about the link between the pentagram map and cross-ratio dynamics.
Remark 7.
As we’ll see in the proof, the infinitesimal monodromy and the polynomial $\chi_{P}$ rewrite themselves as:
$M^{\prime}$ $=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
$\chi_{P}(X,Y)$ $=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
Remark 8.
The dynamic of any closed 3-gon is trivial on the moduli space for any system in $\mathbb{P}^{1}$, since the action of $\mathbb{P}\mathrm{GL}_{2}$ is 3-transitive. So on the initial space, the iteration of $T$ is equivalent to the iteration of a projective transformation.
In some systems (like the flat cross-ratio dynamics [10, Lemma 5.3]), the iteration of $T^{2}$ on closed 4-gon is also trivial on the moduli space. This is reminiscent of classical results on the pentagram map (see [4, §2]). It provides one more evidence in favour of conjecture 3.2.
Proof.
We proceed as in the proof of [7, Prop. 4.1]. By applying the scaling symmetry, we deform $M$ into
$M_{z}=\left({\begin{array}[]{cc}0&c_{1}(z)\\ -1&1\\ \end{array}}\right)\dots\left({\begin{array}[]{cc}0&c_{n}(z)\\ -1&1\\ \end{array}}\right).$
For $1\leq i\leq n$, define $M_{z,i}$ to be the product of the first $i$ matrices (so $M_{z,n}=M_{z}$), and set $M_{z,0}=\mathrm{Id}$. Note that
$\left({\begin{array}[]{cc}0&c_{i}(z)\\ -1&1\\ \end{array}}\right)\dots\left({\begin{array}[]{cc}0&c_{n}(z)\\ -1&1\\ \end{array}}\right)=M_{z,i-1}^{-1}M_{z}.$
This gives that
$\frac{dM_{z}}{dz}=\left(\sum_{i=1}^{n}M_{z,i-1}\left({\begin{array}[]{cc}0&c_{i}^{\prime}(z)\\ 0&0\\ \end{array}}\right)M_{z,i}^{-1}\right)M_{z}.$
Since $M_{1}=\mathrm{Id}$, then the infinitesimal monodromy is
$M^{\prime}:=\left.\frac{dM_{z}}{dz}\right|_{z=1}=\sum_{i=1}^{n}S_{i},\quad\text{where }S_{i}:=M_{1,i-1}\left({\begin{array}[]{cc}0&c_{i}^{\prime}(1)\\ 0&0\\ \end{array}}\right)M_{1,i}^{-1}.$
Since we chose the projective chart where $p_{0}=1,p_{1}=\infty,p_{2}=0$, we obtain the lifts $V_{i}\in k^{2}\setminus\{(0,0)\}$ such that
$V_{0}=\begin{pmatrix}1\\ 1\end{pmatrix},V_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},V_{2}=\begin{pmatrix}0\\ 1\end{pmatrix}$.
By proceeding like in [10, Lemma 3.2], we get that the matrices act this way:
$M_{1,1}:\begin{cases}V_{0}&\mapsto c_{1}V_{1}\\ V_{1}&\mapsto-V_{2}\\ V_{2}&\mapsto V_{3}\\ \begin{pmatrix}c_{2}\\ 1\end{pmatrix}&\mapsto V_{4}\\ \end{cases}.$
Indeed, the first three lines are given by direct computation, and the last one comes from the observation that:
$c_{2}=[1,\infty,0,c_{2}]=[p_{1},p_{2},p_{3},M_{1,1}(c_{2})].$
By definition of $c_{2}$ we obtain that $M_{1,1}(c_{2})=p_{4}$, which means that:
$M_{1,1}\begin{pmatrix}c_{2}\\ 1\end{pmatrix}=V_{4}.$
Similarly, we get by recurrence that for $2\leq i\leq n$:
$M_{1,i}:\begin{cases}V_{0}&\mapsto-c_{i}V_{i}\\ V_{1}&\mapsto-V_{i+1}\\ V_{2}&\mapsto V_{i+2}\\ \begin{pmatrix}c_{i+1}\\ 1\end{pmatrix}&\mapsto V_{i+3}\end{cases}.$
Since $p_{i}\neq p_{i+1}$, then $(V_{i},V_{i+1})$ forms a basis of $k^{2}$. By a simple computation we get:
$S_{i}:\begin{cases}V_{i}&\mapsto\frac{c_{i}^{\prime}(1)}{c_{i}}V_{i}\\ V_{i+1}&\mapsto\begin{pmatrix}0\\ 0\end{pmatrix}\end{cases}$.
We can decompose any $V\in k^{2}$ as
$V=\frac{|V,V_{i+1}|}{|V_{i},V_{i+1}|}V_{i}+\frac{|V_{i},V|}{|V_{i},V_{i+1}|}V_{i+1},$
where $|,|$ denotes the determinant of the $2\times 2$ matrix made out the column vectors. In this basis, the linear map associated to $S_{i}$ is:
$V\mapsto\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{|V,V_{i+1}|}{|V_{i},V_{i+1}|}V_{i}.$
Notice that this linear map commutes with the diagonal action of $\mathrm{GL}_{d+1}$ on the vectors. Indeed, the ratio of the determinants remains unchanged. Moreover, the previous linear map is independent of the choice of lifts $V_{i}$’s. So up to a projective transformation, we can suppose that $p_{i}\neq\infty$ for all $i$ and choose the lifts:
$V_{i}=\begin{pmatrix}p_{i}\\ 1\end{pmatrix}\quad\forall i\in\mathbb{Z}.$
In this choice of affine chart, the matrix expression of the linear map is:
$S_{i}=\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$.
By summing them, we conclude that:
$M^{\prime}=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$.
A small computation makes the quantities $I,J,K$ appear. Note that the eigenvectors of the matrix
$\begin{pmatrix}n/2+J&-K\\ I&n/2-J\end{pmatrix}$
are the same that the ones of
$\begin{pmatrix}J&-K\\ I&-J\end{pmatrix}$.
The eigenvalues change, but it doesn’t matter for our purpose. Let $\begin{pmatrix}X\\ Y\end{pmatrix}$ be an eigenvector with eigenvalue $\lambda$, then:
$\begin{cases}JX-KY&=\lambda X\\ IX-JY&=\lambda Y\end{cases}.$
By multiplying the first line by $Y$ and the second by $X$, we get by subtracting them:
$IX^{2}-2JXY+KY^{2}=0.$
This concludes the proof. ∎
Corollary 4.3.
The quantities $I,J,K$ are invariant by the dynamic. Furthermore, the quantity $\Delta=:J^{2}-IK$ is invariant under the action of $\mathbb{P}\mathrm{GL}_{2}$, but the quantities $I,J,K$ are not. They change under the action of the generators as the following:
$(z\mapsto z-\lambda)\cdot[I,2J,K]$ $=[I,-(2I\lambda-2J),I\lambda^{2}-2J\lambda+K]$
$=[I,-\chi^{\prime}_{P}(\lambda,1),\chi_{P}(\lambda,1)],\qquad\penalty 10000\ \lambda\in k,$
$(z\mapsto\lambda z)\cdot[I,2J,K]$ $=[I\lambda^{-1},2J,K\lambda],\qquad\qquad\qquad\lambda\in k^{*},$
$(z\mapsto\frac{-1}{z})\cdot[I,2J,K]$ $=[K,-2J,I].$
Proof.
The invariance of $I,J,K$ under the dynamic simply comes from the invariance of the infinitesimal monodromy.
The quantity $\Delta=J^{2}-IK$ is the determinant of the polynomial $\chi_{P}$, and so it is projectively invariant.
The action of projective transformations does not change the cross-ratio coordinates, but only the $p_{i}$’s. The modifications of $I,J,K$ under the action of the generators are given by a simple computation. ∎
Some objects keep reappearing in each polygonal dynamic in $\mathbb{P}^{1}$, but we couldn’t make a unifying statement. Let’s conclude this section with open questions.
Open question 1.
Define the following 1-form and (usually presymplectic) 2-form:
$\lambda=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{dp_{i}+dp_{i+1}}{p_{i+1}-p_{i}},\quad\Omega=d\lambda=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{dp_{i}\wedge dp_{i+1}}{(p_{i}-p_{i+1})^{2}}.$
Long specific computations show that in each system, $\Omega$ is invariant through the dynamic, meaning $T^{*}\Omega=\Omega$ (see [10, Thm. 15] and [11, Cor. 5.15]). Could we prove it for any polygonal dynamic in $\mathbb{P}^{1}$?
Moreover, if we consider the generators of the infinitesimal action of $\mathbb{P}\mathrm{GL}_{2}$
$u$ $=\sum_{k\in\mathbb{Z}/n\mathbb{Z}}\frac{\partial}{\partial p_{k}},$ $v$ $=\sum_{k\in\mathbb{Z}/n\mathbb{Z}}p_{k}\frac{\partial}{\partial p_{k}},$ $w=$ $\sum_{k\in\mathbb{Z}/n\mathbb{Z}}p_{k}^{2}\frac{\partial}{\partial p_{k}},$
then a little computation similar to the one done in [10, §6.2] shows that
$i_{u}\Omega$ $=dI,$ $i_{v}\Omega$ $=dJ,$ $i_{w}\Omega$ $=dK+\sum_{i=1}^{n}\frac{1}{2}(\frac{c_{i}^{\prime}(1)}{c_{i}}-\frac{c_{i-1}^{\prime}(1)}{c_{i-1}})dp_{i}.$
If all the $\frac{c_{i}^{\prime}(1)}{c_{i}}$ are equal [^10] For the systems considered here, it only happens for the flat cross-ratio dynamics with constant discrete curvature., then the last sum vanish and this tells us that $I,J,K$ are the hamiltonians of respectively $u,v,w$.
Open question 2.
For $1\leq k\leq\lfloor n/2\rfloor$, consider the quantities
$G_{k}(P)=\sum_{i_{1}<\dots<i_{k}}\frac{c_{i_{1}}^{\prime}(1)}{c_{i_{1}}}\dots\frac{c_{i_{k}}^{\prime}(1)}{c_{i_{k}}}\frac{(p_{i_{1}}-p_{i_{k}+1})(p_{i_{2}}-p_{i_{1}+1})\dots(p_{i_{k}}-p_{i_{k-1}+1})}{(p_{i_{1}}-p_{i_{1}+1})(p_{i_{2}}-p_{i_{2}+1})\dots(p_{i_{k}}-p_{i_{k}+1})}.$
They are projective invariant, since they are sums of multiratios. They are introduced for the flat cross-ratio dynamics in [10, §5.3], where the authors prove the invariance through the dynamic. They also link them to the concept of $\lambda$-length, and to logarithms of alternated perimeters coming from ideal polygons.
A long and painful computation (not reproduced here) shows that these quantities are also invariant for the staircase cross-ratio dynamic. But for the general case of polygonal dynamic over $\mathbb{P}^{1}$, we didn’t manage to prove the invariance. Maybe the answer could come from the $\lambda$-length interpretation, more detailed in [29], which would be coherent from the observation that
$\frac{c_{i}^{\prime}(1)}{c_{i}}=\left.\frac{d\mathrm{ln}(c_{i}(z))}{dz}\right|_{z=1}.$
5 Two examples of dynamics in $\mathbb{P}^{1}$
In this section, we study two polygonal dynamics previously defined in other papers. For any of them, we start by deriving their scaling symmetries, and then apply theorem 4.2 to get their infinitesimal monodromies. From this, we get what are their possible collapse points.
5.1 The leapfrog map
The leapfrog map is defined in [11, §5.2] as an analogue of the pentagram map in $\mathbb{P}^{1}$, and the authors notice that it seems to be the hyperbolic counterpart of the polygon recutting defined in [3] (which would be the euclidean case). For a matter of internal coherence, the notations of [11] are changed to fit the ones from here.
The dynamic involves two $n$-gons
$S^{-}=(s_{i}^{-})_{i\in\mathbb{Z}},\quad S=(s_{i})_{i\in\mathbb{Z}},$
with the same monodromy, which we can intertwine this way
$\dots,s_{i}^{-},s_{i},s_{i+1}^{-},s_{i+1},\dots$
to form a $2n$-gon denoted by $P$. The leapfrog map is defined by
$\Phi:(S^{-},S)\mapsto(S,S^{+}),$
where $S^{+}$ is given by applying local “leapfrog” moves. Such a move maps $s_{i}^{-}$ to $s_{i}^{+}$ via the only projective transformation that fixes $s_{i}$ and interchanges $s_{i-1}$ with $s_{i+1}$. Note that such moves are involutions, and that we could study the group generated by them, like we will do afterwards for the staircase cross-ratio dynamics. However, it is not the point here.
Because of the indexation, it is more convenient work with
$o_{i}^{-}=[s_{i}^{-},s_{i},s_{i-1},s_{i+1}^{-}],\quad o_{i}=[s_{i},s_{i+1}^{-},s_{i}^{-},s_{i+1}],$
instead of the usual $c_{i}$. This gives a fairly simple scaling symmetry.
Proposition 5.1.
There is a scaling symmetry acting by:
$o_{i}^{-}(t)=to_{i}^{-},\qquad o_{i}(t)=o_{i}.$
It is well defined for any $t\in\mathbb{G}_{m}$.
Proof.
First of all, it is direct to check that this is indeed a group action from $\mathbb{G}_{m}$.
We will express the system of coordinates $(o^{-},o)$ with the help of two other systems, and then get the result. According to [11, Prop. 5.12], we have the coordinate systems $(x,y)$ and $(p,r)$, defined by:
$x_{i}$ $=[s_{i+1},s_{i+1}^{-},s_{i+2}^{-},s_{i}^{-}],$ $y_{i}$ $=\frac{(s_{i+1}^{-}-s_{i+1})(s_{i+2}^{-}-s_{i+2})(s_{i}^{-}-s_{i+1}^{-})}{(s_{i+1}^{-}-s_{i+2})(s_{i}^{-}-s_{i+1})(s_{i+1}^{-}-s_{i+2}^{-})},$
$p_{i}$ $=[s_{i+1}^{-},s_{i+2}^{-},s_{i+1},s_{i+2}],$ $r_{i}$ $=[s_{i},s_{i+1},s_{i+1}^{-},s_{i+2}^{-}].$
They are linked by the following relations [^11] The original paper forgot the minus signs, but it doesn’t impact their reasoning.:
$p_{i}$ $=-\frac{y_{i}}{x_{i}},$ $r_{i}$ $=-\frac{x_{i-1}x_{i}}{x_{i-1}(1-x_{i})+y_{i-1}}.$
Now, because of the permutation of the variables of the cross-ratio and the previous relation, we see that:
$o_{i}^{-}$ $=\frac{r_{i-1}}{r_{i-1}-1}=\frac{x_{i-2}x_{i-1}}{x_{i-2}+y_{i-2}},$ $o_{i}$ $=\frac{p_{i-1}}{p_{i-1}-1}=\frac{y_{i-1}}{x_{i-1}+y_{i-1}}.$
But according to [11, Rmk. 3.5], there is a scaling symmetry (actually working for all of their generalised pentagram maps) defined for any $t\in\mathbb{G}_{m}$ by
$x_{i}(t)$ $=tx_{i},$ $y_{i}(t)$ $=ty_{i}.$
Plugging it back to our formula linking $(o^{-},o)$ to $(x,y)$, we get the result. ∎
Now we derive the infinitesimal monodromy and potential collapse points for the leapfrog map.
Proposition 5.2.
The infinitesimal monodromy for the leapfrog map is
$M^{\prime}=\sum_{i=1}^{n}\frac{1}{s_{i}^{-}-s_{i}}\begin{pmatrix}s_{i}^{-}&-s_{i}^{-}s_{i}\\ 1&-s_{i}\end{pmatrix}$,
and its eigenvectors $(X,Y)^{T}$ are roots of the polynomial
$\chi_{P}(X,Y)=\sum_{i=1}^{n}\frac{(X-s_{i}^{-}Y)(X-s_{i}Y)}{s_{i}^{-}-s_{i}}.$
Proof.
We have from proposition 5.1 that
$\frac{o_{i}^{-}\penalty 10000\ {}^{\prime}(1)}{o_{i}^{-}}=1,\qquad\frac{o_{i}^{\prime}(1)}{o_{i}}=0,$
and by remembering that they correspond to the cross-ratio coordinates $c_{i}$, we get the result from applying theorem 4.2. ∎
5.2 Flat cross-ratio dynamics
The flat cross-ratio dynamic was studied in [10] and [19] for constant discrete curvature $\alpha\in\mathbb{C}^{*}$, and for varying $\alpha_{i}\in\mathbb{C}^{*}$ in [16]. Hamiltonian integrability was proved in both situations.
Definition 5.3.
Let $n\geq 3$ and $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon in $\mathbb{P}^{1}$ such that $p_{i}\neq p_{i+1},p_{i+2}$ for each $i\in\mathbb{Z}$. Enrich it with a discrete curvature $\alpha=(\alpha_{1},\dots,\alpha_{n})$ such that $\alpha_{i}\in\mathbb{G}_{m}\setminus\{1\}$ for each $i\in\mathbb{Z}$.
We say that two $n$-gons $P$ and $Q$ are $\alpha$-related if they share the same monodromy, the same discrete curvature, and verify
$\alpha_{i}=[p_{i},p_{i+1},q_{i},q_{i+1}]\quad\forall i\in\mathbb{Z}.$
In a more compact way, we write $P\overset{\alpha}{\sim}Q$.
From [10, Cor. 2.7] we get that for an $n$-gon $P$, there are generically only two $n$-gon $Q^{+},Q^{-}$ that are $\alpha$-related to it. This gives an evolution operator in the futur/past, called the flat cross-ratio dynamics.
Next lemma is an adaptation of [10, Lemma 4.4] for non-constant $\alpha_{i}$’s.
Lemma 5.4.
Let $P$ and $Q$ be two $n$-gons such that $P\overset{\alpha}{\sim}Q$. For all $i\in\mathbb{Z}$, set
$c_{i}$ $=[p_{i},p_{i+1},p_{i-1},p_{i+2}],$ $x_{i}$ $=[p_{i},p_{i+1},p_{i-1},q_{i}],$
$d_{i}$ $=[q_{i},q_{i+1},q_{i-1},q_{i+2}],$ $y_{i}$ $=[q_{i},q_{i+1},q_{i-1},p_{i}].$
Then we have
$\begin{split}c_{i}=\alpha_{i}x_{i}(1-x_{i+1}),\quad d_{i}=\alpha_{i}y_{i}(1-y_{i+1}),\\ \quad 1=x_{i}+y_{i}+x_{i}y_{i}\left(\frac{1-\alpha_{i}}{1-\alpha_{i-1}}-1\right).\end{split}$ (1)
Conversely, for any $(x_{i})\in(\mathbb{G}_{m}\setminus\{1\})^{n}$ and any $(\alpha_{i})\in(\mathbb{G}_{m}\setminus\{1\})^{n}$, the $(c_{i})$ and $(d_{i})$ given by the previous formulas define a pair $P\overset{\alpha}{\sim}Q$, unique up to a simultaneous projective transformation.
Proof.
Again, it is similar to [10]. By using the “Chasles’ relation of the cross-ratio”, we have:
$c_{i}=[p_{i},p_{i+1},p_{i-1},q_{i}][p_{i},p_{i+1},q_{i},q_{i+1}][p_{i},p_{i+1},q_{i+1},p_{i+2}]=x_{i}\alpha_{i}(1-x_{i+1}),$
and similarly we get the same one for $d_{i}$ by exchanging the $p_{i}$’s with $q_{i}$’s, and $x_{i}$’s with $y_{i}$’s. We also have
$[p_{i},q_{i},p_{i-1},p_{i+1}]$ $=[p_{i},q_{i},p_{i-1},q_{i-1}][p_{i},q_{i},q_{i-1},q_{i+1}][p_{i},q_{i},q_{i+1},p_{i+1}]$
$=(1-\alpha_{i-1})[p_{i},q_{i},q_{i-1},q_{i+1}]\frac{1}{1-\alpha_{i}}$
so by referring to the permutations of the variables of the cross-ratio, we get:
$\frac{x_{i}}{1-x_{i}}(1-\alpha_{i})=\frac{1-y_{i}}{y_{i}}(1-\alpha_{i-1}).$
This rewrites as:
$1=x_{i}+y_{i}+x_{i}y_{i}\left(\frac{1-\alpha_{i}}{1-\alpha_{i-1}}-1\right).$
Conversely, the data of $x$ and $\alpha$ allows to reconstruct $y$, $c$ and $d$. We extend them to get $n$-periodic sequences. From $c$, we can choose representant $P=(p_{i})_{i\in\mathbb{Z}}$ such $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$. Because of lemma 4.1, we know its monodromy $M_{P}$. Similarly we construct a polygon $Q=(q_{i})_{i\in\mathbb{Z}}$, with monodromy $M_{Q}$. We observe that
$p_{i},p_{i+1},p_{i-1},q_{i}]=x_{i}=x_{i+n}=[M_{P}(p_{i}),M_{P}(p_{i+1}),M_{P}(p_{i-1}),M_{Q}(q_{i})],$
so by a classical property of the cross-ratio and projective transformations, we necessary have $M_{P}=M_{Q}$. Because of the computation that gave us
$c_{i}=\alpha_{i}x_{i}(1-x_{i+1}),$
we recognize that $p_{i},p_{i+1},q_{i},q_{i+1}]=\alpha_{i}$, so we have indeed $P\overset{\alpha}{\sim}Q$. ∎
This crucial lemma allows us to define a scaling symmetry. This was already observed in [10] for constant $\alpha_{i}$, although the scaling was much simpler and the concept of scaling symmetry was not named.
Proposition 5.5.
We have a scaling symmetry $\rho$ that acts both on the $c_{i}$’s and the $\alpha_{i}$’s, in the following way:
$\alpha_{i}(t)=1+\kappa_{i}(t\alpha_{1}-1),\quad c_{i}(t)=c_{i}\frac{\alpha_{i}(t)}{\alpha_{i}},\qquad\text{where }\kappa_{i}=\frac{1-\alpha_{i}}{1-\alpha_{1}}.$
This is well defined for $t\neq\frac{1}{\alpha_{1}},\frac{\kappa_{i}-1}{\kappa_{i}\alpha_{1}}$.
In any case, this scaling symmetry is well defined around $t=1$. Indeed, $1/\alpha_{1}\neq 1$ and $(\kappa_{i}-1)/\kappa_{i}\alpha_{1}=1$ would imply $\alpha_{i}=0$, which is not the case.
Remark 9.
If all the $\alpha_{i}$ are equal, we get back the scaling symmetry form [10, Lemma 4.4].
Proof.
Let’s check first that $\rho$ is indeed a group action. First, we have
$\alpha_{i}(1)=1+\frac{\alpha_{i}-1}{\alpha_{1}-1}(\alpha_{1}-1)=\alpha_{i},$
and so $c_{i}(1)=c_{i}$. Furthermore, we note that
$\kappa_{i}(t)=\frac{1-\alpha_{i}(t)}{1-\alpha_{1}(t)}=\frac{\kappa_{i}(t\alpha_{1}-1)}{(t\alpha_{1}-1)}=\kappa_{i},$
so $\kappa_{i}(t)$ is constant for each $i$ ($\rho$ was in fact designed from that property). Then we have
$\alpha_{i}(t_{1})(t_{2})$ $=1+\kappa_{i}(t_{2}\alpha_{1}(t_{1})-1)$
$=1+\kappa_{i}(t_{2}t_{1}\alpha_{1}-1)$
$=\alpha_{i}(t_{1}t_{2}),$
and from this we obtain
$c_{i}(t_{1})(t_{2})$ $=c_{i}(t_{1})\frac{\alpha_{i}(t_{1})(t_{2})}{\alpha_{i}(t_{1})}$
$=c_{i}\frac{\alpha_{i}(t_{1})}{\alpha_{i}}\frac{\alpha_{i}(t_{1}t_{2})}{\alpha_{i}(t_{1})}$
$=c_{i}(t_{1}t_{2}).$
So $\rho$ is indeed a group action. It is well defined if $\alpha_{i}(t)\neq 0,1$ for each $i$, which corresponds to $t\neq\frac{1}{\alpha_{1}},\frac{\kappa_{i}-1}{\kappa_{i}\alpha_{1}}$.
Let’s now see why it commutes with the evolution operator. Take $c,d,x,y$ as in lemma 5.4. They are linked by equation 1. Note that $\frac{1-\alpha_{i}}{1-\alpha_{i-1}}$ is equal to $\frac{\kappa_{i}}{\kappa_{i-1}}$, so it stays constant when applying $\rho$. In fact, applying $\rho$ modifies $c,d,\alpha$, but leaves $x$ and $y$ unchanged. The equation 1 still holds for $c_{i}(t),d_{i}(t),\alpha_{i}(t),x_{i},y_{i}$ and so by lemma 5.4 we have $P(t)\overset{\alpha(t)}{\sim}Q(t)$. Hence $\rho$ is indeed a scaling symmetry. ∎
Proposition 5.6.
The infinitesimal monodromy for the flat cross-ratio dynamics is
$M^{\prime}=\sum_{i=1}^{n}\frac{(1-\alpha_{i})\alpha_{1}}{(1-\alpha_{1})\alpha_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
and its eigenvectors $(X,Y)^{T}$ are roots of the polynomial
$\chi_{P}(X,Y)=\sum_{i=1}^{n}\frac{(1-\alpha_{i})}{\alpha_{i}}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
Proof.
From proposition 5.5 we see that
$\frac{c_{i}^{\prime}(1)}{c_{i}}=\frac{(1-\alpha_{i})\alpha_{1}}{(1-\alpha_{1})\alpha_{i}},$
so by theorem 4.2 we get the formula for the infinitesimal monodromy. Still from the same theorem, we obtain that an eigenvector $(X,Y)^{T}$ is a root of the polynomial $\chi_{P}(X,Y)$. Up to multiplying it by $\frac{\alpha_{1}}{(1-\alpha_{1})}$, which we do only for cosmetic purpose, we get desired formula for the polynomial. ∎
6 A group dynamic in $\mathbb{P}^{1}$: staircase cross-ratio dynamics
We define a new polygonal dynamic in $\mathbb{P}^{1}$ given by a group of polygonal transformations. As said in the introduction, it is a generalization of other systems coming from the continuation of discrete holomorphic functions (see [12, §6] and references therein).
6.1 Definition and results
Before defining the group, let’s start with a few preliminaries.
Definition 6.1.
Let $\alpha,\beta\in\mathbb{G}_{m}$. Two couples of distinct points $(a,c),(b,d)$ of $\mathbb{P}^{1}$ form a $\bm{\alpha/\beta}$-quadrangle when $a,c,b,d]=\alpha/\beta$. The point $c$ is called the $\bm{\alpha/\beta}$-conjugate of $a$ with respect to $(b,d)$, and we write $c=h^{\alpha/\beta}_{b,d}(a)$. When $\alpha/\beta=-1$, the quadrangle said to be harmonic.
Proposition 6.2.
The following statements are true.
  1. 1.
    Projective transformations send $\alpha/\beta$-quadrangles to $\alpha/\beta$-quadrangles. Moreover, we have:
    $g\circ h^{\alpha/\beta}_{b,d}=h^{\alpha/\beta}_{g(b),g(d)}\circ g,\quad\text{for all }g\in\mathbb{P}\mathrm{GL}_{2}.$
  2. 2.
    The $\alpha/\beta$-conjugation with respect to $(b,d)$ is a projective transformation $h^{\alpha/\beta}_{b,d}(z)=\frac{(\beta d-\alpha b)z-(\beta-\alpha)bd}{(\beta-\alpha)z-(\beta b-\alpha d)}$, which fixes $b,d,$ and has characteristic constant $\alpha/\beta$. Its inverse is $h^{\alpha/\beta}_{d,b}$.
  3. 3.
    If $(b,d)$ is fixed, then for $\alpha,\beta,\gamma,\delta\in\mathbb{G}_{m}$, we have $h^{\alpha/\beta}_{b,d}\circ h^{\gamma/\delta}_{b,d}=h^{\alpha\gamma/\beta\delta}_{b,d}$. Hence $h^{1}_{b,d}=id$ and $(h^{\alpha/\beta}_{b,d})^{-1}=h^{\beta/\alpha}_{b,d}$.
Proof.
Let’s check each statement.
  1. 1.
    Let $(a,c),(b,d)$ be an $\alpha/\beta$-quadrangle and $g\in\mathbb{P}\mathrm{GL}_{2}$. The cross-ratio is $\mathbb{P}\mathrm{GL}_{2}$-invariant, so
    $g(a),g(h^{\alpha/\beta}_{b,d}(a)),g(b),g(d)]=\frac{\alpha}{\beta}=[g(a),h^{\alpha/\beta}_{g(b),g(d)}(g(a)),g(b),g(d)],$
    which gives the result.
  2. 2.
    The formula of $h^{\alpha/\beta}_{b,d}$ is the result of direct calculation. To see the characteristic constant, it suffices to conjugate by a projective transformation that sends $(b,d)$ to $(\infty,0)$, use (a) and by the formula we see that $h^{\alpha/\beta}_{\infty,0}:z\mapsto\frac{\alpha}{\beta}z$.
    Considering $(c,a),(d,b)$ also gives a $\alpha/\beta$-quadrangle, because this permutation doesn’t modify the cross-ratio. So we get
    $a=h^{\alpha/\beta}_{d,b}(c)=h^{\alpha/\beta}_{d,b}\circ h^{\alpha/\beta}_{b,d}(a),$
    hence $h^{\alpha/\beta}_{d,b}=(h^{\alpha/\beta}_{b,d})^{-1}$.
  3. 3.
    Let $g\in\mathbb{P}\mathrm{GL}_{2}$ which sends $(b,d)$ to $(\infty,0)$. Using the (a) and (b), we obtain
    $h^{\alpha/\beta}_{b,d}\circ h^{\gamma/\delta}_{b,d}$ $=g^{-1}\circ h^{\alpha/\beta}_{\infty,0}\circ g\circ g^{-1}\circ h^{\gamma/\delta}_{\infty,0}\circ g$
    $=g^{-1}\circ h^{\alpha\gamma/\beta\delta}_{\infty,0}\circ g$
    $=h^{\alpha\gamma/\beta\delta}_{b,d}.\qed$
We can now define the dynamic.
Definition 6.3.
Let $n\geq 3$ and $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon in $\mathbb{P}^{1}$ such that $p_{i}\neq p_{i+1},p_{i+2}$ for each $i\in\mathbb{Z}$. Enrich it with an $n$-periodic discrete curvature $\mu=(\mu_{i})_{i\in\mathbb{Z}}$ such that $\mu_{i}\in\mathbb{G}_{m}$ for each $i\in\mathbb{Z}$.
Let $j\in\mathbb{Z}/n\mathbb{Z}$. The flip at index $\bm{j}$ is the map
[…]
such that:
$\begin{cases}\tilde{p}_{i}=h^{\mu_{i}/\mu_{i-1}}_{p_{i+1},p_{i-1}}(p_{i})&\quad\forall i\in n\mathbb{Z}+j,\\ \tilde{p}_{i}=p_{i}&\quad\forall i\notin n\mathbb{Z}+j,\\ \tilde{\mu}_{i-1}=\mu_{i}&\quad\forall i\in n\mathbb{Z}+(j-1),\\ \tilde{\mu}_{i}=\mu_{i-1}&\quad\forall i\in n\mathbb{Z}+j,\\ \tilde{\mu}_{i}=\mu_{i}&\quad\forall i\notin n\mathbb{Z}+(j-1)\cup n\mathbb{Z}+j.\end{cases}$
Because of proposition 6.2(1), $\tilde{P}$ has the same monodromy as $P$. In a more compact way, we write $P\overset{\mu}{\sim}_{j}\tilde{P}$. See figure 2 for a visualisation of the action of $\phi_{j}$.
Define the group of flips to be $\Gamma_{n}:=\langle\phi_{1},\dots,\phi_{n}\rangle$. Its dynamic is called the staircase cross-ratio dynamics.
Remark 10.
The dynamic remains the same if we consider the discrete curvature $\lambda\mu:=(\lambda\mu_{i})_{i\in\mathbb{Z}}$ for $\lambda\in\mathbb{G}_{m}$, since we’re only interested in ratios.
Remark 11.
This dynamical system consists of local moves, contrarily to the flat cross-ratio dynamic. Because of this, it is also well-defined if we add a monodromy $\gamma\in\mathbb{G}_{m}$ on the discrete curvature $\mu$:
$\mu_{i+n}=\gamma\mu_{i},\quad\forall i\in\mathbb{Z}.$
Moreover, one could consider the case of discrete curves: these are polygons $(p_{i})_{i\in\mathbb{Z}}$ endowed with discrete curvature $(\mu_{i})_{i\in\mathbb{Z}}$ which are not satisfying any monodromy conditions. We won’t study them here, but it might be an interesting subject.
[…]
Figure 2: Visualisation of the action of the flip $\phi_{j}$ at index $j$, read from up to down. It it similar to a braid on a cylinder, but it is important not to forget that it acts also on the polygon $(p_{i})_{i\in\mathbb{Z}/n\mathbb{Z}}$; indeed, the point $p_{j}$ is changed in $\tilde{p}_{j}=h^{\mu_{j}/\mu_{j-1}}_{p_{j+1},p_{j-1}}(p_{j})$.
Proposition 6.4.
Let $\Gamma_{n}=\langle\phi_{1},\dots,\phi_{n}\rangle$ be the group generated by flips. It has presentation
$\langle\phi_{1},\dots,\phi_{n}\mid\phi_{i}^{2}=1,\quad\phi_{i}\phi_{j}=\phi_{j}\phi_{i}\text{ if }|i-j|\geq 2,\quad\phi_{i}\phi_{i+1}\phi_{i}=\phi_{i+1}\phi_{i}\phi_{i+1}\rangle,$
where the indices (and the distance) are understood modulo $n$. This is usually called the affine symmetric group (see [13]).
Proof.
As it was done by Adler in [3] for the polygon recutting, we will check each relations.
  • $\phi_{i}^{2}=1$: The action of $\phi_{i}$ switches $\mu_{i-1}$ with $\mu_{i}$, and applies $h^{\mu_{i}/\mu_{i-1}}_{p_{i+1},p_{i-1}}$ on $p_{i}$. Applying it again switches back $\mu_{i}$ with $\mu_{i-1}$, and apply $h^{\mu_{i-1}/\mu_{i}}_{p_{i+1},p_{i-1}}$ on $h^{\mu_{i}/\mu_{i-1}}_{p_{i+1},p_{i-1}}(p_{i})$. So the $\mu_{i-1},\mu_{i}$ are back at their place, and with the proposition 6.2(3) we have $h^{\mu_{i-1}/\mu_{i}}_{p_{i+1},p_{i-1}}\circ h^{\mu_{i}/\mu_{i-1}}_{p_{i+1},p_{i-1}}(p_{i})=p_{i}$. Hence the result.
  • $\phi_{i}\phi_{j}=\phi_{j}\phi_{i}\text{ if }|i-j|\geq 2$: This comes directly from the fact that $\phi_{i}$ only acts on $\mu_{i-1},\mu_{i}$ and $p_{i}$ (using $p_{i-1},p_{i+1}$). So any change on indices $j$ with $|i-j|\geq 2$ doesn’t affect anything on how $\phi_{i}$ will act, and we have the commutation.
  • $\phi_{i}\phi_{i+1}\phi_{i}=\phi_{i+1}\phi_{i}\phi_{i+1}$: This comes from the reinterpretation of a property called “3D-consistency”([30, Prop. 9]). It states that if the vertices $(p_{i-1},p_{i},p_{i+1},p_{i+2})$ and edges $(\mu_{i-1},\mu_{i},\mu_{i+1})$ are positioned on a cube like on figure 3, one can “flip around the cube” by doing $(\phi_{i+1}\phi_{i})^{3}$ and get back to the same state. So $(\phi_{i+1}\phi_{i})^{3}=\mathrm{Id}$, and since flips are involutions we get the result.∎
$p_{i}$$p_{i-1}$$p_{i+1}$$p_{i+2}$$\tilde{p}_{i}$$\tilde{p}_{i+1}$$\tilde{\tilde{p}}_{i}$$\tilde{\tilde{p}}_{i+1}$$\mu_{i}$$\mu_{i}$$\mu_{i}$$\mu_{i}$$\mu_{i+1}$$\mu_{i+1}$$\mu_{i+1}$$\mu_{i+1}$$\mu_{i-1}$$\mu_{i-1}$$\mu_{i-1}$$\mu_{i-1}$
Figure 3: A visualization of 3D consistency, meaning that one can “flip around the cube consistently”. The red vertices and edges are the ones taken originally.
Remark 12.
The group $\Gamma_{n}$ doesn’t always act faithfully. It is the case for instance when the field $k$ is finite (since any orbit is finite). However, when there is an element $q\in k$ of infinite multiplicative order [^12] For instance $2\in\mathbb{Q}$., then each iterate of $\psi:=\phi_{n}\circ\dots\circ\phi_{1}\in\Gamma_{n}$ acts differently (see subsection 6.2). Hence the dynamic doesn’t consist only of finites orbits, which is coherent with our conjectures.
Next lemma is the analogue of lemma 5.4. It characterises the dynamic on the moduli space, for any flip $\phi_{j}$. Notice that it holds for discrete curves.
Lemma 6.5.
Let $P=((p_{1},\dots,p_{n}),M)$ be an $n$-gon with discrete curvature $\mu=(\mu_{1},\dots,\mu_{n})$. Let $j\in\mathbb{Z}$ and $\tilde{P}=\phi_{j}(P)$. For all $i\in\mathbb{Z}$, set
$c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}],\quad\tilde{c}_{i}=[\tilde{p}_{i},\tilde{p}_{i+1},\tilde{p}_{i-1},\tilde{p}_{i+2}],$
and set
$x_{j}=[p_{j-2},p_{j-1},\tilde{p}_{j},p_{j}],\quad\tilde{x}_{j}=[\tilde{p}_{j-2},\tilde{p}_{j-1},p_{j},\tilde{p}_{j}],$
$y_{j}=[p_{j+1},p_{j+2},p_{j},\tilde{p}_{j}],\quad\tilde{y}_{j}=[\tilde{p}_{j+1},\tilde{p}_{j+2},\tilde{p}_{j},p_{j}],$
such that
$x_{j}\tilde{x}_{j}=1=y_{j}\tilde{y}_{j}.$
Then we have
$c_{i}=\tilde{c}_{i}\quad\forall i\neq j-2,j-1,j,j+1,$ ($\star_{i}$)
$c_{j-2}=\tilde{c}_{j-2}x_{j},$ ($\star_{j-2}$)
$c_{j-1}(\tilde{c}_{j-1}-1)\mu_{j-1}=\tilde{c}_{j-1}(c_{j-1}-1)\tilde{\mu}_{j-1},$ ($\star_{j-1}$)
$c_{j}(\tilde{c}_{j}-1)\mu_{j}=\tilde{c}_{j}(c_{j}-1)\tilde{\mu}_{j},$ ($\star_{j}$)
$c_{j+1}=\tilde{c}_{j+1}y_{j}.$ ($\star_{j+1}$)
Conversely, for any fixed $j\in\mathbb{Z}$, $x_{j},y_{j}\in\mathbb{G}_{m}$ and $(\mu_{i})\in(\mathbb{G}_{m})^{n}$, any $(c_{i})$ and $(\tilde{c}_{i})$ verifying the previous equations define a pair $P\overset{\mu}{\sim}_{j}\tilde{P}$, unique up to a simultaneous projective transformation.
Proof.
This proof is quite similar to the one of lemma 5.4. Since the flip $\phi_{j}$ leaves the vertices $p_{i}$ untouched for $i\notin n\mathbb{Z}+j$, we have that
$c_{i}=\tilde{c}_{i}\quad\forall i\neq j-2,j-1,j,j+1.$
Hence $\star_{i}$ is proved. We also obtain that $x_{j}=\tilde{x}_{j}^{-1}$ and $y_{j}=\tilde{y}_{j}^{-1}$.
By the “Chasles’ relation” of the cross ratio, we have:
$c_{j-2}$ $=[p_{j-2},p_{j-1},p_{j-3},p_{j}]$
$=[p_{j-2},p_{j-1},p_{j-3},\tilde{p}_{j}][p_{j-2},p_{j-1},\tilde{p}_{j},p_{j}]$
$=\tilde{c}_{j-2}x_{j},$
$\frac{\mu_{j-1}}{\tilde{\mu}_{j-1}}$ $=[\tilde{p}_{j},p_{j},p_{j+1},p_{j-1}]$
$=[\tilde{p}_{j},p_{j-2},p_{j+1},p_{j-1}][p_{j-2},p_{j},p_{j+1},p_{j-1}]$
$=\frac{\tilde{c}_{j-1}}{\tilde{c}_{j-1}-1}\frac{c_{j-1}-1}{c_{j-1}},$
$\frac{\mu_{j}}{\tilde{\mu}_{j}}$ $=[p_{j},\tilde{p}_{j},p_{j+1},p_{j-1}]$
$=[p_{j},p_{j+2},p_{j+1},p_{j-1}][p_{j+2},\tilde{p}_{j},p_{j+1},p_{j-1}]$
$=\frac{c_{j}-1}{c_{j}}\frac{\tilde{c}_{j}}{\tilde{c}_{j}-1},$
$c_{j+1}$ $=[p_{j+1},p_{j+2},p_{j},p_{j+3}]$
$=[p_{j+1},p_{j+2},p_{j},\tilde{p}_{j}][p_{j+1},p_{j+2},\tilde{p}_{j},p_{j+3}]$
$=y_{j}\tilde{c}_{j+1}.$
Hence we get $\star_{j-2}$, $\star_{j-1}$, $\star_{j}$, $\star_{j+1}$.
Conversely, choose some representants $P=(p_{i})_{i\in\mathbb{Z}}$ and $\tilde{P}=(\tilde{p}_{i})_{i\in\mathbb{Z}}$, equipped with discrete curvatures $\mu$ and $\tilde{\mu}$, such that $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$ and $\tilde{c}_{i}=[\tilde{p}_{i},\tilde{p}_{i+1},\tilde{p}_{i-1},\tilde{p}_{i+2}]$. Because of lemma 4.1, we know their monodromies $M_{P}$, $M_{\tilde{P}}$.
The equations $\star_{i}$ give us equalities between many pairs of cross-ratios. Hence there exists a projective transformation $g$ such that:
$\tilde{p}_{i}=g(p_{i})\quad\forall i\in\{j+1,\dots,j+n-1\}.$
We have:
$c_{j+1}$ $=[p_{j+1},p_{j+2},p_{j},p_{j+3}]$
$=[p_{j+1},p_{j+2},p_{j},\tilde{p}_{j}][p_{j+1},p_{j+2},\tilde{p}_{j},p_{j+3}]$
$=y_{j}[g(p_{j+1}),g(p_{j+2}),g(p_{j}),g(\tilde{p}_{j+3})]$
$=y_{j}[\tilde{p}_{j+1},\tilde{p}_{j+2},\tilde{p}_{j},g(\tilde{p}_{j+3})].$
By $\star_{j+1}$ we obtain that
$\tilde{p}_{j+1},\tilde{p}_{j+2},\tilde{p}_{j},g(\tilde{p}_{j+3})]=\tilde{c}_{j+1}=[\tilde{p}_{j+1},\tilde{p}_{j+2},\tilde{p}_{j},\tilde{p}_{j+3}],$
which forces $g$ to be the identity, so
$\tilde{p}_{i}=p_{i}\quad\forall i\in\{j+1,\dots,j+n-1\}.$
By doing almost the same and using $\star_{j-2}$ instead, we obtain that:
$\tilde{p}_{i}=p_{i}\quad\forall i\in\{j-n+1,\dots,j-1\}.$
This gives us $M_{P}=M_{\tilde{P}}$, because they both map $(p_{j-n+1},p_{j-n+2},p_{j-n+3})$ to $(p_{j+1},p_{j+2},p_{j+3})$. Hence we have that:
$p_{i}=\tilde{p}_{i}\quad\forall i\notin n\mathbb{Z}+j.$
By rearranging the equation $\star_{j}$ in the same way as in the beginning of this proof, we see that indeed:
$\tilde{p}_{j}=h^{\mu_{j}/\mu_{j-1}}_{p_{j+1},p_{j-1}}(p_{j}).$
Hence $P\overset{\mu}{\sim}_{j}\tilde{P}$, which concludes the proof. ∎
Now we can use the previous lemma to get a family of scaling symmetries. They don’t depend on the index $j$ where the flip is performed. Notice that again, they still hold for discrete curves. However, they are not adapted to $n$-gons with a monodromy on $\mu$.
Proposition 6.6.
For each $\eta\in\mathbb{G}_{m}$, there is a scaling symmetry $\rho^{(\eta)}$ that acts in the following way:
$\mu_{i}^{(\eta)}(t)=\frac{t\mu_{i}}{1+(t-1)\eta\mu_{i}},\quad c_{i}^{(\eta)}(t)=c_{i}\frac{t\mu_{i}}{\mu_{i}^{(\eta)}(t)}.$
This is well defined for $t\neq 1-(\eta\mu_{i})^{-1}$.
Remark 13.
Considering the scaling symmetries for each $\eta\in\mathbb{G}_{m}$ is useful. Indeed, if we consider the discrete curvature $\lambda\mu$ for $\lambda\in\mathbb{G}_{m}$, then we have the following commutative diagram.
$(c,\mu)$$(c^{(\eta)}(t),\mu^{(\eta)}(t))$$(c,\lambda\mu)$$(c^{(\lambda\eta)}(t),\lambda\mu^{(\lambda\eta)}(t))$$\scriptstyle{\rho^{(\eta)}(t)}$$\scriptstyle{\lambda}$$\scriptstyle{\lambda}$$\scriptstyle{\rho^{(\lambda\eta)}(t)}$
However, we’ll see later that the eigenlines of the infinitesimal monodromy don’t depend on the choice of $\eta$.
Proof.
Like for the proof of proposition 5.5, we have first to check that $\rho^{(\eta)}$ is indeed a group action. First, we clearly have $\mu_{i}^{(\eta)}(1)=\mu_{i}$ and $c_{i}^{(\eta)}(1)=c_{i}$. Then we have
$\mu_{i}^{(\eta)}(t_{1})(t_{2})$ $=\frac{t_{2}\mu_{i}^{(\eta)}(t_{1})}{1+(t_{2}-1)\eta\mu_{i}^{(\eta)}(t_{1})}$
$=\frac{t_{2}t_{1}\mu_{i}}{1+(t_{1}-1)\eta\mu_{i}}\frac{1}{1+(t_{2}-1)\frac{t_{1}\mu_{i}}{1+(t_{1}-1)\eta\mu_{i}}}$
$=\frac{t_{1}t_{2}\mu_{i}}{1+(t_{1}-1)\eta\mu_{i}+t_{1}(t_{2}-1)\eta\mu_{i}}$
$=\frac{t_{1}t_{2}\mu_{i}}{1+(t_{1}t_{2}-1)\eta\mu_{i}}$
$=\mu_{i}^{(\eta)}(t_{1}t_{2}),$
and from this we get
$c_{i}^{(\eta)}(t_{1})(t_{2})$ $=c_{i}^{(\eta)}(t_{1})\frac{t_{2}\mu_{i}^{(\eta)}(t_{1})}{\mu_{i}^{(\eta)}(t_{1})(t_{2})}$
$=c_{i}\frac{t_{1}\mu_{i}}{\mu_{i}^{(\eta)}(t_{1})}\frac{t_{2}\mu_{i}^{(\eta)}(t_{1})}{\mu_{i}^{(\eta)}(t_{1}t_{2})}$
$=c_{i}\frac{t_{1}t_{2}\mu_{i}}{\mu_{i}^{(\eta)}(t_{1}t_{2})}$
$=c_{i}^{(\eta)}(t_{1}t_{2}).$
So $\rho^{(\eta)}$ acts indeed as a group. It is well defined when every $\mu_{i}^{(\eta)}(t)$ is in $\mathbb{G}_{m}$, which means that $t\neq 1-(\eta\mu_{i})^{-1}$.
Then, because of the converse part of lemma 6.5, we need to check that equations $\star_{i}$, $\star_{j-2}$, $\star_{j-1}$, $\star_{j}$ and $\star_{j+1}$ hold for the pairs $(c^{(\eta)}(t),\mu^{(\eta)}(t))$ and $(\tilde{c}^{(\eta)}(t),\tilde{\mu}^{(\eta)}(t))$.
First of all, we have that $c_{i}^{(\eta)}(t)=\tilde{c}_{i}^{(\eta)}(t)$ for each $i\neq j-2,j-1,j,j+1$, so $\star_{i}$ is still verified.
Then, we have
$x_{i}\tilde{c}_{j-2}^{(\eta)}(t)=x_{i}\tilde{c}_{j-2}\frac{t\mu_{j-2}}{\mu_{j-2}^{(\eta)}(t)}=c_{j-2}\frac{t\mu_{j-2}}{\mu_{j-2}^{(\eta)}(t)}=c_{j-2}^{(\eta)}(t),$
so $\star_{j-2}$ is still verified. We get $\star_{j+1}$ in the exact same way. Furthermore, we have
$c_{j}^{(\eta)}(t)(\tilde{c}_{j}^{(\eta)}(t)-1)\mu_{j}^{(\eta)}(t)$ $=c_{j}\frac{t\mu_{j}}{\mu_{j}^{(\eta)}(t)}(\tilde{c}_{j}\frac{t\tilde{\mu_{j}}}{\tilde{\mu}_{j}^{(\eta)}(t)}-1)\mu_{j}^{(\eta)}(t)$
$=c_{j}t\mu_{j}(\tilde{c_{j}}(1+(t-1)\eta\tilde{\mu}_{j})-1)$
$=tc_{j}(\tilde{c}_{j}-1)\mu_{j}+t(t-1)c_{j}\tilde{c}_{j}\eta\mu_{j}\tilde{\mu}_{j}$
$=t\tilde{c}_{j}(c_{j}-1)\tilde{\mu}_{j}+t(t-1)\tilde{c}_{j}c_{j}\eta\tilde{\mu}_{j}\mu_{j}$
$=\tilde{c}_{j}t\tilde{\mu}_{j}(c_{j}(1+(t-1)\eta\mu_{j})-1)$
$=\tilde{c}_{j}^{(\eta)}(t)(c_{j}^{(\eta)}(t)-1)\tilde{\mu}_{j}^{(\eta)}(t),$
so $\star_{j}$ is still verified, and the same computation also gives us $\star_{j-1}$. ∎
Now we can compute the infinitesimal monodromy and the coordinate of the collapse points [^13] The polynomial for the collapse points was first found experimentally by Paul Melotti, when working on discrete holomorphic functions.. Figure 4 depicts the convergence of one randomly chosen closed $5$-gon, under the iteration of some non-trivial element of $\Gamma_{5}$, towards one of the collapse points, as predicted by conjecture 3.2.
Proposition 6.7.
The infinitesimal monodromy for the staircase cross-ratio dynamics, under the scaling symmetry $\rho^{(\eta)}$, is
$M^{\prime}=\sum_{i=1}^{n}\eta\mu_{i}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
and its eigenvectors $(X,Y)^{T}$ are roots of the polynomial
$\chi_{P}(X,Y)=\sum_{i=1}^{n}\mu_{i}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
Remark 14.
If we define $\alpha_{i}:=[p_{i},p_{i+1},p_{i-1},\tilde{p}_{i}]$ to match the convention of the flat cross-ratio dynamics, we have by the permutation of the variables of the cross-ratio that
$\mu_{i}=[p_{i},\tilde{p}_{i},p_{i+1},p_{i-1}]=\frac{\alpha_{i}-1}{\alpha_{i}}.$
So, up to an unimportant sign, we get back the same formula as in proposition 5.6. This could be expectable, since they both arise from the theory of discrete holomorphic functions.
Proof.
From proposition 6.6 we have that
$\frac{c_{i}^{\prime}(1)}{c_{i}}=\frac{c_{i}\eta\mu_{i}}{c_{i}}=\eta\mu_{i},$
and we get the infinitesimal monodromy by theorem 4.2. We also get the polynomial $\chi_{P}$, which is of the desired form when we divide it by $\eta$. ∎
Figure 4: Python simulation of the evolution of one randomly chosen closed $5$-gon in $\mathbb{P}^{1}(\mathbb{C})$ with randomly chosen discrete curvature, under the iteration of the sequence of flips $\phi_{5}\circ\dots\circ\phi_{1}$ (applied 25 times). The initial vertices are represented with red crosses. A new edge is drawn each time a flip is done, and the colour indicates time (from blue to red). There is a rapid spiraling towards one of the collapse points determined by corollary 3.4.
6.2 Special configurations
[…]
Figure 5: Special “staircase” of the cross-ratio dynamic, for $n=3$. On the left, this makes appear a parabolic Möbius transformation, and on the right an elliptic or loxodromic (depending on $|q|$).
Here, we investigate some specific cases of staircase cross-ratio dynamics for which we can write explicitly the dynamic. Their motion on the moduli space is periodic, meaning that it amounts on the initial space to iterate a projective transformation (of any kind). They can be seen as a special case of discrete holomorphic functions, depicted in figure 5. The idea comes from [31], and it holds on any algebraicly closed field of characteristic 0. For an example of generic orbit in $\mathbb{P}^{1}(\mathbb{C)}$, see the simulation in figure 4.
First, let’s see that how a parabolic transformation can appear. Let’s consider a closed $n$-gon
$P=(0,1,\dots,n-1)$
with discrete curvature $(1,\dots,1,\lambda)$, such that after applying $\phi_{n}\circ\dots\circ\phi_{1}$, we get
$\tilde{P}=(n,n+1,\dots,2n-1).$
This happens if the cross-ratio on any square is:
$\lambda=\frac{(k-(k+n-1)(k+n-(k+1))}{(k+n-1-(k+n))(k+1-k)}=(n-1)^{2}.$
Hence, we get the Möbius transformation $z\mapsto z+n$, which is parabolic. The dynamic converges in the future and past towards the double point $\infty$.
Now, let’s do the same for a closed $n$-gon
$P=(1,q,\dots,q^{n-1})$
with discrete curvature $(1,\dots,1,\lambda)$, such that after applying $\phi_{n}\circ\dots\circ\phi_{1}$ we get
$\tilde{P}=(q^{n},q^{n+1},\dots,q^{2n-1}).$
Again, we have to check the cross-ratio condition for every square:
$\lambda=\frac{(q^{k}-q^{k+n-1})(q^{k+n}-q^{k+1})}{(q^{k+n-1}-q^{k+n})(q^{k+1}-q^{k})}=\frac{(1-q^{n-1})(q^{n}-q)}{(q^{n-1}-q^{n})(q-1)}.$
By remembering that $\frac{q^{n-1}-1}{q-1}=\sum_{k=0}^{n-2}q^{k}$ and doing some simplifications, we get that $q$ is a root of the polynomial equation
$Q_{n}(q,\lambda)=(\sum_{k=0}^{n-2}q^{k})^{2}-\lambda q^{n-2}=0.$
The polynomial $Q_{n}$ is of degree 2n42n-4 and palindromic with respect to $q$, which implies that $1/q$ is also a root. So we can suppose $|q|\geq 1$. Note that if $q=1$, we get back $\lambda=(n-1)^{2}$ as in the parabolic case.
Let’s go back to conjecture 3.2. In the case of periodic motion of the moduli space, any type of Möbius transformation can appear.
  1. 1.
    If $|q|\neq 1$, we get a convergence towards $\infty$ in the future and 0 in the past (loxodromic transformation, the generic case);
  2. 2.
    If $q$ is a root of unity, then the dynamic is periodic (finite order elliptic transformation);
  3. 3.
    If $|q|=1$ but is not a root of unity, then the dynamic is recurrent (infinite order elliptic transformation);
  4. 4.
    Otherwise, we get a convergence in the past and future towards $\infty$ (parabolic transformation).
In all of this cases, we can deform the polygons via the scaling symmetry to get families of twisted $n$-gons. The generic case is the first, which is coherent with conjectures 3.2, but the other ones remind us that it won’t hold everywhere.
Funding
This work was partially funded by the French National Research Agency under the project DynAtrois (ANR-24-CE40-1163)
Acknowledgements
I warmly thank my PhD advisors Thomas Gauthier and Ramanujan Santharoubane for their constant support, and also Paul Melotti with whom I started working on this question during my master’s. He was the first to have the intuition of the collapse formula for the system that then became the staircase cross-ratio dynamics. I’d also like to thank Niklas Affolter, Anton Izosimov, Sanjay Ramassamy, Richard Schwartz, Valentin Ovsienko and Max Weinreich for the helpful and sympathetic discussions.
nocite*